OpenStudy (anonymous):

A firm had retained earnings of $100,000 in 2009 and $125,000 in 2010. The increase in retained earnings from 2009 to 2010 is: A. 12.5 percent. B. 20 percent. C. 25 percent. D. 125 percent.

OpenStudy (anonymous):

I might be doing my math wrong but isn't it C?

OpenStudy (anonymous):

why 25?

OpenStudy (anonymous):

Well math is 100,000 to 125,000 is 25%? Because fractions: 1/4 to 2/4 is 25%

OpenStudy (anonymous):

Wait I'm looking at that now it's A xD

OpenStudy (anonymous):

Durp

OpenStudy (anonymous):

Sorry about that :/

OpenStudy (anonymous):

i wound say 125 percent but not sure 100000/100=1000 then 125000/100=1250 1000/1250=1.25*100=125

OpenStudy (anonymous):

You have it almost right but your forgetting the decimal. The most logical answer is A It can't be 125% because then he would have to go from $0 to $125,000!

OpenStudy (anonymous):

the answer is 12.5?

OpenStudy (anonymous):

It has to be but if you want to wait for someone to prove it be my guest :) I'm alright with math and I'm thinking that 12.5 has to be the answer.

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

No problem

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."